# Eigenvector decorrelation for random matrices

### v2

@@ 4.3.3. Conditional Gronwall estimates: regular case
Next, for $m\in\{k-1,k\}$, we note that (by simple combinatorics) there are at least two $R$’s, which are hit by a derivative exactly once. Therefore, using (6.12) for all the terms originating from the other $m-2$ differentiated $R$’s, and ignoring the “untouched" $|R|^{2p-m}$ factor, we arrive at
@@
(6.15)$\prec\left(\frac{1}{N^{k/2}\eta\gamma^{2}}+\frac{\mathcal{E}_{0}}{N^{(k-1)/2}\eta\gamma}+\frac{\mathcal{E}_{0}^{2}}{N^{(k-2)/2}\eta}\right)\,\big{(}\gamma^{-1}+\mathcal{E}_{0}\big{)}^{m-2}\lesssim\left(1+\frac{1}{\sqrt{N}\eta^{7/6}}\right)\mathcal{E}_{1}^{m}\,.$
@@
$N^{-k/2}\sum_{l=0}^{k}\left|\sum_{a,b}\mathbf{E}(\partial_{ab}^{l}\partial_{ba}^{k-l}|R|^{2p})\right|\lesssim\left(1+\frac{1}{\sqrt{N}\eta^{3/2}}\right)\left(\mathbf{E}|R|^{2p}+N^{\xi}\mathcal{E}_{1}^{2p}\right)\,,$
@@
(6.16) $\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{E}|R_{t}|^{2p}\lesssim\sum_{k=3}^{K}\frac{1}{N^{k/2}}\sum_{l=0}^{k}\left|\sum_{a,b}\mathbf{E}(\partial_{ab}^{l}\partial_{ba}^{k-l}|R_{t}|^{2p})\right|+\mathcal{O}(N^{-100p})\,.$
@@
$N^{-5/2}\left|\sum_{ab}(G_{1}B_{1}G_{2}B_{2}G_{1})_{ab}M_{aa}M_{bb}\right|\lesssim N^{-5/2}\left|(G_{1}B_{1}G_{2}B_{2}G_{1})_{\bm{m}\bm{m}}\right|\prec\frac{1}{N^{3/2}\eta\gamma}\lesssim\mathcal{E}_{1}$
@@
(6.20) $\begin{split}&N^{-9/2}\left|\sum_{a,b}\big{(}(G_{1}B_{1}G_{2}B_{2}G_{1})_{ab}\big{)}^{3}\right|\\ \prec\,&N^{-7/2}\frac{1}{\eta\gamma}\max_{a}\,(G_{1}B_{1}G_{2}B_{2}G_{1}G_{1}^{*}B_{2}^{*}G_{2}^{*}B_{1}^{*}G_{1}^{*})_{aa}\\ \lesssim\,&\frac{1}{N^{7/2}\eta_{1}^{3}\eta_{2}^{2}}\max_{a}\,(G_{1}B_{1}\Im G_{2}B_{1}^{*}G_{1}^{*})_{aa}\lesssim\frac{1}{\sqrt{N}\eta}\,\frac{1}{(N\eta_{1}\eta_{2})^{3}}=\frac{\mathcal{E}_{1}^{3}}{\sqrt{N}\eta}\,.\end{split}$
@@
$\begin{split}N^{-3/2}&\left|\sum_{a,b}G_{\bm{x}a}(G_{1}B_{1}G_{2})_{b\bm{y}}M_{\bm{x}a}G_{bb}(G_{1}B_{1}G_{2})_{a\bm{y}}\right|\\ &\prec N^{-3/2}\frac{1}{\eta}\sum_{a}|G_{\bm{x}a}|^{2}\sum_{b}|(G_{1}B_{1}G_{2})_{b\bm{y}}|\prec\frac{1}{N\eta^{3}}=\mathcal{E}_{0}^{2}\end{split}$
@@ Appendix A Proofs of additional technical results
In this appendix, we collect several results of technical results, that were used in the main text.
@@
[…]
@@
[…]
@@
Assume for simplicity that $\mathbf{I}_{1}=\mathbf{I}_{2}=\mathbf{R}$. Since $\lVert D_{j}\rVert\leq L$, we have that ${\rm supp}\rho_{j}\subset[-L-2,L+2]$ for $j=1,2$. In the following, we will distinguish the two cases (i) $\max\{|z_{1}|,|z_{2}|\}\geq L+3$ and (ii) $\max\{|z_{1}|,|z_{2}|\}\leq L+3$.
@@
$d_{1}:={\rm dist}(z_{1},{\rm supp}\rho_{1})=\min\{|z_{1}-x|:x\in{\rm supp}\rho_{1}\}.$
@@
$\langle M_{1}M_{1}^{*}\rangle=\frac{\langle\Im M_{1}\rangle}{\eta_{1}+\langle\Im M_{1}\rangle}\leq\frac{1}{1+d^{2}}.$
@@
$1\gtrsim\beta_{*}(z_{1},z_{2})\geq 1-\max\{|\langle M_{1}M_{2}\rangle|,|\langle M_{1}M_{2}^{*}\rangle\}\geq 1-\langle M_{1}M_{1}^{*}\rangle^{1/2}\langle M_{2}M_{2}^{*}\rangle^{1/2}\geq\frac{d^{2}}{1+d^{2}}\gtrsim 1$
Here we used that $\langle M_{2}M_{2}^{*}\rangle\leq 1$ and $d\geq 1$. Moreover, $\eta_{1}/\rho_{1}\gtrsim d^{2}$, which implies $\widehat{\gamma}(z_{1},z_{2})\sim 1$. Thus $\beta_{*}(z_{1},z_{2})\sim\widehat{\gamma}(z_{1},z_{2})$.
@@
Lower bound on $\beta_{*}$. Taking into account [48, Proposition 4.2] it is sufficient to show that $\mathrm{LT}\lesssim\beta_{*}$. Subtracting (1.4) for $M_{1}$ from (1.4) for $M_{2}^{*}$ we get that
@@
In the last step we used (3.4) which is proven in [48, Proposition 4.2]. For the second factor we use the bound
@@
Vague monotonicity in space: Note that $\widehat{\gamma}$ has the symmetry $\widehat{\gamma}(z_{1},z_{2},D_{1},D_{2})=\widehat{\gamma}(z_{2},z_{1},D_{2},D_{1})$. Thus it is sufficient to prove the first part of (4.18). In the following, we will distinguish between the two cases (i) $|\langle M_{1}(z_{1})M_{2}^{*}(z_{2})\rangle|\geq 1/2$ and (ii) $|\langle M_{1}(z_{1})M_{2}^{*}(z_{2})\rangle|<1/2$. The exact choice of the threshold separating this two cases is not important, $1/2$ may be replaced by any $c\in(0,1)$. The proof in case (ii) is much simpler, since it corresponds to the situation when $\beta_{*}(z_{1},z_{2})\gtrsim 1$ and one only needs to show that $\beta_{*}(z_{1},w_{2})\gtrsim 1$. The proof in case (i), however, is much more involved.
@@
After having treated the case $|\langle M_{1}(z_{1})M_{2}^{*}(w_{2})\rangle|\geq 1/2$, we may assume that $|\langle M_{1}(z_{1})M_{2}^{*}(w_{2})\rangle|<1/2$. In this case, we have $\beta(z_{1},\bar{w}_{2})\geq 1/2$. Notice that
@@
If $\rho_{2}(w_{2})<1/(8L)$, then (A.17) gives $\beta(z_{1},w_{2})\geq 1/4$. Otherwise by [48, Proposition 4.2] it holds that $\beta(z_{1},w_{2})\gtrsim\rho_{2}(w_{2})^{2}\gtrsim 1$. This means that $\beta_{*}(z_{1},w_{2})\sim 1$. Therefore, by Proposition 3.1 $\gamma_{0}(z_{1},w_{2})\sim 1$, which immediately implies (A.11).
@@
which immediately follows from Lemma 3.3 applied to $y_{1}=y_{2}=0$. This finishes the verification of (4.22b). $\square$
@@
(A.22) $M_{1}(z_{1}^{\prime})-M_{2}(z_{2})=\frac{(z_{1}^{\prime}-z_{2})-\langle M_{1}(z_{1}^{\prime})(D_{1}-D_{2})M_{2}(z_{2})\rangle}{1-\langle M_{1}(z_{1}^{\prime})M_{2}(z_{2})\rangle}M_{1}(z_{1}^{\prime})M_{2}(z_{2})-M_{1}(z_{1}^{\prime})(D_{1}-D_{2})M_{2}(z_{2}).$
@@
$|\langle M_{2}(z_{2}^{\prime})A(M_{1}(z_{1}^{\prime})-M_{2}(z_{2}))^{*}\rangle|\lesssim\lVert A\rVert\left(|z_{1}^{\prime}-z_{2}|+\langle(D_{1}-D_{2})^{2}\rangle^{1/2}\right)\lesssim\lVert A\rVert\widehat{\gamma}(z_{1}^{\prime},z_{2})\lesssim\lVert A\rVert\widehat{\gamma}(z_{1}^{\prime},z_{2}^{\prime}).$
@@
$\int_{s}^{t}f_{r}\mathrm{d}r\leq 2\int_{s}^{t}|\langle M_{12,r}^{I}\rangle|\mathrm{d}r\leq\int_{s}^{r}\left(\frac{\pi\rho_{1,r}}{\eta_{1,r}}+\frac{\pi\rho_{2,r}}{\eta_{2,r}}\right)\mathrm{d}r\leq\log\frac{\eta_{1,s}\eta_{2,s}}{\eta_{1,t}\eta_{2,t}}.$
@@
$\beta_{s}=|1-\langle M_{1,s}M_{2,s}\rangle|=|1-\mathrm{e}^{s-t}\langle M_{1,t}M_{2,t}\rangle|=|\mathrm{e}^{s-t}(1-\langle M_{1,t}M_{2,t}\rangle)+(1-\mathrm{e}^{s-t})|\sim\beta_{t}+t-s,$
@@
In the last step we used that $(x-E)^{2}+\eta^{2}\sim E^{2}+\eta^{2}$ for any $x\in{\rm supp}\,\rho_{T}$ once the distance from $E$ to the support of $\rho$ has a lower bound of order 1. Therefore it holds that
@@
$[0,\infty)\ni\eta\mapsto\eta\rho_{T}(E+\mathrm{i}\eta)=\frac{1}{\pi}\int_{\mathbf{R}}\frac{\eta^{2}}{(x-E)^{2}+\eta^{2}}\rho_{T}(x)\mathrm{d}x$
for any fixed $E$. Moreover, since this function increases from 00 at $\eta=0$ to $1$ at $\eta\to+\infty$, for any $E\in\mathbf{R}$ there exists a unique $\eta=\eta(E)>0$ such that
@@
While the rhs. of (A.29) is strictly smaller than $-1$, the lhs. equals
@@
the lhs. of (A.29) is lower bounded by $-1$. Hence, we arrive at a contraction and hence Lemma 5.4 (i) holds.
@@
where $M_{T}$ is evaluated at $z_{1,T}+x+\mathrm{i}x$. We view (A.31) as a lower bound on $T-t$ and are hence left to show that the lhs. of (A.31) has an upper bound of order one.

### v3

@@ 4.3.3. Conditional Gronwall estimates: regular case
Next, for $m\in\{k-1,k\}$, we note that (by simple combinatorics) there are at least two $R$’s, which are hit by a derivative exactly once. Therefore, using (6.10) for all the terms originating from the other m2m-2 differentiated $R$’s, and ignoring the “untouched" $|R|^{2p-m}$ factor, we arrive at
@@
[…]
@@
$N^{-k/2}\sum_{a,b}\sum_{l=0}^{k}\left|\mathbf{E}(\partial_{ab}^{l}\partial_{ba}^{k-l}|R|^{2p})\right|\lesssim\left(1+\frac{1}{\sqrt{N}\eta^{3/2}}\right)\left(\mathbf{E}|R|^{2p}+N^{\xi}\mathcal{E}_{1}^{2p}\right)\,$,
@@
(6.14) $\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{E}|R_{t}|^{2p}\lesssim\sum_{k=3}^{K}\frac{1}{N^{k/2}}\sum_{a,b}\sum_{l=0}^{k}\left|\mathbf{E}(\partial_{ab}^{l}\partial_{ba}^{k-l}|R_{t}|^{2p})\right|+\mathcal{O}(N^{-100p})\,.$
@@
$N^{-5/2}\Big|\sum_{ab}(G_{1}B_{1}G_{2}B_{2}G_{1})_{ab}M_{aa}M_{bb}\Big|\lesssim N^{-5/2}\left|(G_{1}B_{1}G_{2}B_{2}G_{1})_{\bm{m}\bm{m}}\right|\prec\frac{1}{N^{3/2}\eta\gamma}\lesssim\mathcal{E}_{1}$
@@
(6.18) $\begin{split}&N^{-9/2}\Big|\sum_{a,b}\big((G_{1}B_{1}G_{2}B_{2}G_{1})_{ab}\big)^{3}\Big|\\ \prec\,&N^{-7/2}\frac{1}{\eta\gamma}\max_{a}\,(G_{1}B_{1}G_{2}B_{2}G_{1}G_{1}^{*}B_{2}^{*}G_{2}^{*}B_{1}^{*}G_{1}^{*})_{aa}\\ \lesssim\,&\frac{1}{N^{7/2}\eta_{1}^{3}\eta_{2}^{2}}\max_{a}\,(G_{1}B_{1}\Im G_{2}B_{1}^{*}G_{1}^{*})_{aa}\lesssim\frac{1}{\sqrt{N}\eta}\,\frac{1}{(N\eta_{1}\eta_{2})^{3}}=\frac{\mathcal{E}_{1}^{3}}{\sqrt{N}\eta}\,.\end{split}$
@@
$\begin{split}N^{-3/2}&\Big|\sum_{a,b}G_{\bm{x}a}(G_{1}B_{1}G_{2})_{b\bm{y}}M_{\bm{x}a}G_{bb}(G_{1}B_{1}G_{2})_{a\bm{y}}\Big|\\ &\prec N^{-3/2}\frac{1}{\eta}\sum_{a}|G_{\bm{x}a}|^{2}\sum_{b}|(G_{1}B_{1}G_{2})_{b\bm{y}}|\prec\frac{1}{N\eta^{3}}=\mathcal{E}_{0}^{2}\end{split}$
@@ Appendix A Proofs of additional technical results
(A.1) $\beta_{*}(E_{1}+\mathrm{i}0,E_{2}+\mathrm{i}0)\sim\widehat{\gamma}(E_{1}+\mathrm{i}0,E_{2}+\mathrm{i}0).$
@@
In the regime $\Delta^{2}+|E_{1}-E_{2}|\leq c$, the estimate (A.1) follows from a straightforward perturbative calculation for $\beta(E_{1}+\mathrm{i}0,E_{2}-\mathrm{i}0)$. Here, the implicit constant $c>0$ depends only on $\kappa$ and $L$. In the complementary regime, we have $\widehat{\gamma}\sim 1$ and also $\beta_{*}\sim 1$ by Proposition 3.1. Therefore, it holds that $\beta_{*}\sim\widehat{\gamma}$. The rest of the proof of Proposition A.1 is elementary and thus omitted. ∎
@@
Assume for simplicity that $\mathbf{I}_{1}=\mathbf{I}_{2}=\mathbf{R}$. Since $\lVert D_{j}\rVert\leq L$, we have that ${\mathrm{s}upp}\rho_{j}\subset[-L-2,L+2]$ for $j=1,2$. In the following, we will distinguish the two cases (i) $\max\{|z_{1}|,|z_{2}|\}\geq L+3$ and (ii) $\max\{|z_{1}|,|z_{2}|\}\leq L+3$.
@@
$d_{1}:={\mathrm{d}ist}(z_{1},{\mathrm{s}upp}\rho_{1})=\min\{|z_{1}-x|:x\in{\mathrm{s}upp}\rho_{1}\}.$
@@
$\langle M_{1}M_{1}^{*}\rangle=\frac{\langle\Im M_{1}\rangle}{\eta_{1}+\langle\Im M_{1}\rangle}\leq\frac{1}{1+d^{2}_{1}}.$
@@
$\begin{split}1&\gtrsim\beta_{*}(z_{1},z_{2})\geq 1-\max\{|\langle M_{1}M_{2}\rangle|,|\langle M_{1}M_{2}^{*}\rangle\}\\ &\geq 1-\langle M_{1}M_{1}^{*}\rangle^{1/2}\langle M_{2}M_{2}^{*}\rangle^{1/2}\geq 1-\frac{1}{\sqrt{1+d^{2}_{1}}}\gtrsim 1.\end{split}$
Here we used that $\langle M_{2}M_{2}^{*}\rangle\leq 1$ and $d_{1}\geq 1$. Moreover, $\eta_{1}/\rho_{1}\gtrsim d^{2}_{1}$, which implies $\widehat{\gamma}(z_{1},z_{2})\sim 1$. Thus $\beta_{*}(z_{1},z_{2})\sim\widehat{\gamma}(z_{1},z_{2})$.
@@
Lower bound on $\beta_{*}$. Taking into account [49, Proposition 4.2] it is sufficient to show that $\mathrm{LT}\lesssim\beta_{*}$. Subtracting (1.4) for $M_{1}$ from (1.4) for $M_{2}^{*}$ we get that
@@
In the last step we used (3.4) which is proven in [49, Proposition 4.2]. For the second factor we use the bound
@@
Vague monotonicity in space: Note that $\widehat{\gamma}$ has the symmetry
$\widehat{\gamma}(z_{1},z_{2},D_{1},D_{2})=\widehat{\gamma}(z_{2},z_{1},D_{2},D_{1}).$
Thus it is sufficient to prove the first part of (4.18). In the following, we will distinguish between the two cases (i) $|\langle M_{1}(z_{1})M_{2}^{*}(z_{2})\rangle|\geq 1/2$ and (ii) $|\langle M_{1}(z_{1})M_{2}^{*}(z_{2})\rangle|<1/2$. The exact choice of the threshold separating this two cases is not important, $1/2$ may be replaced by any $c\in(0,1)$. The proof in case (ii) is much simpler, since it corresponds to the situation when $\beta_{*}(z_{1},z_{2})\gtrsim 1$ and one only needs to show that $\beta_{*}(z_{1},w_{2})\gtrsim 1$. The proof in case (i), however, is much more involved.
@@
After having treated the case $|\langle M_{1}(z_{1})M_{2}^{*}(w_{2})\rangle|\geq 1/2$, we may assume that
$|\langle M_{1}(z_{1})M_{2}^{*}(w_{2})\rangle|<1/2.$
In this case, we have $\beta(z_{1},\bar{w}_{2})\geq 1/2$. Notice that
@@
If $\rho_{2}(w_{2})<1/(8L)$, then (A.17) gives $\beta(z_{1},w_{2})\geq 1/4$. Otherwise by [49, Proposition 4.2] it holds that $\beta(z_{1},w_{2})\gtrsim\rho_{2}(w_{2})^{2}\gtrsim 1$. This means that $\beta_{*}(z_{1},w_{2})\sim 1$. Therefore, by Proposition 3.1 $\gamma_{0}(z_{1},w_{2})\sim 1$, which immediately implies (A.11).
@@
which immediately follows from Lemma 3.3 applied to $y_{1}=y_{2}=0$. This finishes the verification of (4.22b). ∎
@@
(A.22) $\begin{split}M_{1}(z_{1}^{\prime})-M_{2}(z_{2})&=\frac{(z_{1}^{\prime}-z_{2})-\langle M_{1}(z_{1}^{\prime})(D_{1}-D_{2})M_{2}(z_{2})\rangle}{1-\langle M_{1}(z_{1}^{\prime})M_{2}(z_{2})\rangle}M_{1}(z_{1}^{\prime})M_{2}(z_{2})\\ &-M_{1}(z_{1}^{\prime})(D_{1}-D_{2})M_{2}(z_{2}).\end{split}$
@@
$\begin{split}|\langle M_{2}(z_{2}^{\prime})A(M_{1}(z_{1}^{\prime})-M_{2}(z_{2}))^{*}\rangle|&\lesssim\lVert A\rVert\left(|z_{1}^{\prime}-z_{2}|+\langle(D_{1}-D_{2})^{2}\rangle^{1/2}\right)\\ &\lesssim\lVert A\rVert\widehat{\gamma}(z_{1}^{\prime},z_{2})\lesssim\lVert A\rVert\widehat{\gamma}(z_{1}^{\prime},z_{2}^{\prime}).\end{split}$
@@
$\int_{s}^{t}f_{r}\mathrm{d}r\leq 2\int_{s}^{t}|\langle M_{12,r}^{I}\rangle|\mathrm{d}r\leq\int_{s}^{t}\left(\frac{\pi\rho_{1,r}}{\eta_{1,r}}+\frac{\pi\rho_{2,r}}{\eta_{2,r}}\right)\mathrm{d}r\leq\log\frac{\eta_{1,s}\eta_{2,s}}{\eta_{1,t}\eta_{2,t}}.$
@@
$\begin{split}\beta_{s}&=|1-\langle M_{1,s}M_{2,s}\rangle|=|1-\mathrm{e}^{s-t}\langle M_{1,t}M_{2,t}\rangle|\\ &=|\mathrm{e}^{s-t}(1-\langle M_{1,t}M_{2,t}\rangle)+(1-\mathrm{e}^{s-t})|\sim\beta_{t}+t-s,\end{split}$
@@
In the last step we used that $(x-E)^{2}+\eta^{2}\sim E^{2}+\eta^{2}$ for any $x\in{\mathrm{s}upp}\,\rho_{T}$ once the distance from $E$ to the support of $\rho$ has a lower bound of order 1. Therefore it holds that
@@
$0,\infty)\ni\eta\mapsto\eta\rho_{T}(E+\mathrm{i}\eta)=\frac{1}{\pi}\int_{\mathbf{R}}\frac{\eta^{2}}{(x-E)^{2}+\eta^{2}}\rho_{T}(x)\mathrm{d}x$
for any fixed $E$. Moreover, since this function increases from 0 at $\eta=0$ to $1$ at $\eta\to+\infty$, for any $E\in\mathbf{R}$ there exists a unique $\eta=\eta(E)>0$ such that
@@
While the rhs. of (A.29) is strictly smaller than 1-1, the lhs. equals
@@
the lhs. of (A.29) is lower bounded by 1-1. Hence, we arrive at a contraction and hence Lemma 5.4 (i) holds.
@@
where $M_{T}$ is evaluated at $z_{1,T}+x+\mathrm{i}x$. We view (A.31) as a lower bound on TtT-t and are hence left to show that the lhs. of (A.31) has an upper bound of order one.